# Quasinormal modes of slowly-spinning horizonless compact objects

### v2

@@ Abstract
One of the main predictions of general relativity is the existence of black holes featuring a horizon beyond which nothing can escape. Gravitational waves from the remnants of compact binary coalescences have the potential to probe new physics close to the BH horizons. This prospect is of particular interest given several quantum-gravity models that predict the presence of horizonless and singularity-free compact objects. The membrane paradigm is a generic framework that allows one to parametrise the interior of compact objects in terms of the properties of a fictitious fluid located at the object’s radius. It has been used to derive the quasinormal mode spectrum of static horizonless compact objects. Extending the membrane paradigm to rotating objects is crucial to constrain the properties of the spinning merger remnants. In this work, we extend the membrane paradigm to linear order in spin and use it to analyse the relationships between the quasi-normal modes, the object’s reflectivity, and the membrane parameters. We find a breaking of isospectrality between axial and polar modes when the object is partially reflecting or the compactness differs from the black hole case. We also find that in reflective ultracompact objects some of the modes tend towards instability as the spin increases. Finally, we show that the spin enhances the deviations from the black-hole quasinormal mode spectrum as the compactness decreases. This implies that spinning horizonless compact objects may be more easily differentiated than nonspinning ones in the prompt ringdown.
@@ II.1 Background metric
Motivated by the fact that the remnants of compact binary coalescences have large values of the spin Abbott et al. (2019a, 2021a, 2024, 2021b), here we extend the membrane paradigm to spinning horizonless compact objects which are generally described by the Kerr metric at the first order in the spin. It is worth mentioning that the Kerr geometry does not necessarily describe the vacuum region outside a spinning object due to the absence of Birkhoff’s theorem beyond spherical symmetry. Here, we assume that GR works sufficiently well on the exterior of the compact object, and no specific model is considered for the object’s interior. This assumption is valid in modified theories of gravity where the corrections to the metric are suppressed by powers of $\ell_{P}/r_{0}\ll 1$, being $r_{0}$ the object’s radius and $\ell_{P}$ the Planck length or the scale of new physics Berti et al. (2015).
@@
where $f(r)=1-2M/r$, $d\Sigma^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$, $J=aM$ is the angular momentum, $\chi=a/M$ is the dimensionless spin, and $M$ is the total mass of the compact object. Note that for BHs, the magnitude of the spin parameter is $|\chi|\leq 1$ to prevent naked singularities, but no such constraint exists for generic compact objects. A generic spinning body is also characterized by the spin-induced multipole moments, which are relevant from the second order in spin Vines and Steinhoff (2018); Saketh and Vines (2022). For the Kerr metric, all the multipole moments are known as unique functions of $M$ and $J$. In this work, we truncate to linear order in spin and thus remove contribution of spin-induced multipole moments to the metric. Moreover, the exterior spacetime has no ergoregion, i.e., the region outside the horizon where a static observer cannot exist and is forced to corotate with the compact object.
@@
where $\epsilon>0$ and lies in the range $\mathcal{O}(10^{-40})-\mathcal{O}(1)$ for compact objects with Planckian corrections at the horizon scale and whose compactness is comparable to that of a neutron star, respectively.
@@ II.2 Linear perturbations
where $g_{\mu\nu}^{0}$ is the metric in Eq. (4) and $\epsilon_{p}$ is a small parameter controlling the strength of perturbation. The metric perturbation $\delta g_{\mu\nu}$ can be decomposed in the spherical harmonic basis as Regge and Wheeler (1957); Zerilli (1970)
$\delta g_{\mu\nu}=\delta g_{\mu\nu}^{\rm{axial}}+\delta g_{\mu\nu}^{\rm{polar}}\,,$ (9)
where under parity transformations ($\theta\rightarrow\pi-\theta$, $\phi\rightarrow\phi+\pi$), the axial sector transforms as $(-1)^{\ell+1}$ and the polar sector as $(-1)^{\ell}$. In the absence of spin, the two sectors evolve independently without mixing. However, in the presence of spin, they mix in the field equations and the boundary conditions. In the slow-rotation limit, it can be shown that the coupling terms between the axial and polar sectors can be dropped for the computation of th QNM frequencies to linear order in spin Pani (2013). The axial and polar perturbations are given by Kojima (1992); Pani (2013)
$\delta g_{\mu\nu}^{\text{axial}}$ $=2\sum_{\ell m}\big{\{}h_{0}(r,t)\left[S_{\theta}(\theta,\phi)d\theta+S_{\phi}(\theta,\phi)d\phi\right]dt$
$+h_{1}(r,t)\left[S_{\theta}(\theta,\phi)d\theta+S_{\phi}(\theta,\phi)d\phi\right]dr\big{\}}\,,$ (10)
$\delta g_{\mu\nu}^{\text{polar}}=$ $\sum_{\ell m}\left[H_{0}(r,t)dt^{2}+H_{2}(r,t)dr^{2}+r^{2}K(r,t)d\Sigma^{2}\right.$
$\left.+2H_{1}(r,t)dtdr\right]Y^{\ell m}(\theta,\phi)\,,$ (11)
where $S_{\theta}(\theta,\phi)=-\partial_{\phi}Y_{\ell m}(\theta,\phi)/\sin\theta$, $S_{\phi}(\theta,\phi)=\sin\theta\partial_{\theta}Y_{\ell m}(\theta,\phi)$, $Y_{\ell m}(\theta,\phi)$ are the tensor spherical harmonics, and $\ell$ and $m$ are the angular and azimuthal numbers of the perturbation. The time dependence can be factorised as
@@
where the same holds for $h_{1}(r,t)$, $H_{0}(r,t)$, $H_{1}(r,t)$, $H_{2}(r,t)$, $K(r,t)$, and the explicit dependence of the radial functions on $(\ell,m)$ is dropped for simplicity of notation.
@@ III.1 Reflectivity of compact objects
For generic compact objects, the reflectivity $|\mathcal{R}|$ is related to the membrane parameters (i.e., shear and bulk viscosity of the fluid) through the boundary conditions in Eqs. (II.3) and (II.3) for axial and polar perturbations, respectively. The computation of the reflectivity for arbitrary frequencies is done via numerical integration of the perturbation equations in Eq. (14) in Sec III.1. However, some features of the reflectivity may be extracted analytically and are discussed in the next few paragraphs. We will also discuss the limitations of the linear-in-spin approximation in the form of spurious effects, specifically superradiance, leading to $|\mathcal{R}|>1$ when the spin parameter $\chi$ becomes large enough. This is physically not allowed as superradiance requires the presence of an ergoregion, which is absent at the linear order in spin. Nevertheless, this spuriousness is understood in terms of the limits of validity of the linear-in-spin approximation.
@@ III.1.2 Numerical computation of the reflectivity : Results
Fig. 2 shows the absolute value of the reflectivity of a compact object with $\epsilon\ll 1$ and different values of the shear viscosity, particularly $\eta\in(0,\eta_{\rm{BH}})$ in the top panel and $\eta\in(\eta_{\rm{BH}},+\infty)$ in the bottom panel. We have set $\chi=0.1$ for simplicity. The plots are for axial perturbations, but they look qualitatively similar in the polar case, yielding to the same results. The reflectivity profile coincides with the BH one when $\eta=\eta_{\rm{BH}}$, and describes a perfectly reflecting object in the $\eta\to 0$ and $\eta\to+\infty$ limits. It is interesting to note that for generic values of the shear viscosity, the reflectivity shows an oscillatory structure that becomes resonant when $\eta\to 0$ and $\eta\to+\infty$. This behaviour is due to the QNMs of the compact object, where the frequencies at which the dips (or resonances) occur coincide with the real part of the QNMs of the object. Using the analytical formulae for the QNMs of ultracompact objects in the small-frequency regime Maggio et al. (2019b, 2021b), we check that the distance between subsequent dips (or resonances) is given by
@@
Let us now analyse the variation of the reflectivity with the object’s compactness for a fixed viscosity. For this purpose, we fix the viscosity in the axial boundary condition in Eq. (II.3) to the BH value. The purely ingoing boundary condition is recovered for $\epsilon\rightarrow 0$, as seen from Eq. (32). However, in the limit $\epsilon\rightarrow 1/2$, corresponding to the membrane radius located at the light ring, $r_{0}\rightarrow 3M$, the reflectivity becomes unity as the axial boundary condition tends to the Dirichlet boundary condition, i.e., $\psi_{\rm{RW}}(r_{0})=0$. This happens regardless of the spin as the second and third terms in the right hand side of Eq. (32) are divergent. The top panel of Fig. 3 shows the variation of the reflectivity with the object’s compactness for axial modes in the slowly spinning case where $\chi=0.1$. The reflectivity coincides with the BH one when $\eta=\eta_{\rm{BH}}$, and describes a perfectly reflecting object in the $\epsilon\to 1/2$ limit.
@@
In the polar case, interestingly, the boundary condition in the spinless case ($\chi=0$) does not reach the perfectly reflecting limit as $\epsilon\rightarrow 1/2$. Moreover, the reflectivity depends on the choice of the bulk viscosity of the membrane fluid. However, the linear-in-spin term $H(r_{0},\omega,\eta,\zeta)$ in Eq. (II.3) is divergent in the limit $\epsilon\rightarrow 1/2$, thus changing the behaviour even for a small value of spin. The bottom panel of Fig. 3 shows the variation of the reflectivity with the object’s compactness for polar modes in the slowly spinning case where $\chi=0.1$. Note that the small-spin approximation breaks down for axial and polar modes in the limit $\epsilon\rightarrow 1/2$. This is because the linear-in-spin term of the boundary condition scales as $(\epsilon-1/2)^{-2}$, whereas the spinless term of the boundary condition scales as $(\epsilon-1/2)^{-1}$, as shown in Eq. (II.3). Nevertheless, due to the boundary condition diverging in the limit $\epsilon\rightarrow 1/2$ both in the nonspinning and spinning cases, for $\epsilon=1/2$ the compact object is purely reflecting.

### v3

@@ Abstract
One of the main predictions of general relativity is the existence of black holes featuring a horizon beyond which nothing can escape. Gravitational waves from the remnants of compact binary coalescences have the potential to probe new physics close to the black hole horizons. This prospect is of particular interest given several quantum-gravity models that predict the presence of horizonless and singularity-free compact objects. The membrane paradigm is a generic framework that allows one to parametrise the interior of compact objects in terms of the properties of a fictitious fluid located at the object’s radius. It has been used to derive the quasinormal mode spectrum of static horizonless compact objects. Extending the membrane paradigm to rotating objects is crucial to constrain the properties of the spinning merger remnants. In this work, we extend the membrane paradigm to linear order in spin and use it to analyse the relationships between the quasi-normal modes, the object’s reflectivity, and the membrane parameters. We find a breaking of isospectrality between axial and polar modes when the object is partially reflecting or the compactness differs from the black hole case. We also find that in reflective ultracompact objects some of the modes tend towards instability as the spin increases. Finally, we show that the spin enhances the deviations from the black-hole quasinormal mode spectrum as the compactness decreases. This implies that spinning horizonless compact objects may be more easily differentiated than nonspinning ones in the prompt ringdown.
@@ II.1 Background metric
Motivated by the fact that the remnants of compact binary coalescences have large values of the spin Abbott et al. (2019a, 2021a, 2024, 2021b), here we extend the membrane paradigm to spinning horizonless compact objects which are generally described by the Kerr metric at the first order in the spin. Our approach should be considered as a first step towards the inclusion of spin effects into compact bodies, as performed in literature for compact objects beyond Kerr and in modified theories of gravity Pani et al. (2012, 2013); Uchikata and Yoshida (2016); Pierini and Gualtieri (2021); Wagle et al. (2022). The linear-in-spin approximation is also useful for understanding the new features and conceptual challenges that arise with the inclusion of spin. The extension to higher orders in the spin is left for future work following Refs. Hartle and Thorne (1968); Pierini and Gualtieri (2022); Wagle et al. (2024) It is worth mentioning that the Kerr geometry does not necessarily describe the vacuum region outside a spinning object due to the absence of Birkhoff’s theorem beyond spherical symmetry. Here, we assume that GR works sufficiently well on the exterior of the compact object, and no specific model is considered for the object’s interior. This assumption is valid in modified theories of gravity where the corrections to the metric are suppressed by powers of $\ell_{P}/r_{0}\ll 1$, being $r_{0}$ the object’s radius and $\ell_{P}$ the Planck length or the scale of new physics Berti et al. (2015).
@@
where $f(r)=1-2M/r$, $d\Sigma^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$, $J=aM$ is the angular momentum, and $M$ is the total mass of the compact object. For future convenience, we also define $\chi=a/M$ as the dimensionless spin. Note that for BHs, the magnitude of the spin parameter is $|\chi|\leq 1$ to prevent naked singularities, but no such constraint exists for generic compact objects. A generic spinning body is also characterized by the spin-induced multipole moments, which are relevant from the second order in spin Vines and Steinhoff (2018); Saketh and Vines (2022). For the Kerr metric, all the multipole moments are known as unique functions of $M$ and $J$. In this work, we truncate to linear order in spin and thus remove contribution of spin-induced multipole moments to the metric. Moreover, the exterior spacetime has no ergoregion, i.e., the region outside the horizon where a static observer cannot exist and is forced to corotate with the compact object.
@@
where $\epsilon>0$, particularly $\epsilon=\mathcal{O}(10^{-40})$ for compact objects with Planckian corrections at the horizon scale, and $\epsilon=\mathcal{O}(1)$ for objects whose compactness is comparable to that of a neutron star.
@@ II.2 Linear perturbations
where $g_{\mu\nu}^{0}$ is the metric in Eq. (4) and $\epsilon_{p}$ is a small parameter controlling the strength of perturbation. The metric perturbation $\delta g_{\mu\nu}$ can be split into two parity sectors as Regge and Wheeler (1957); Zerilli (1970)
$\delta g_{\mu\nu}=\delta g_{\mu\nu}^{\rm{axial}}+\delta g_{\mu\nu}^{\rm{polar}}\,.$ (9)
In the absence of spin, the two sectors evolve independently without mixing. However, in the presence of spin, they mix in the field equations and the boundary conditions. In the slow-rotation limit, it can be shown that the coupling terms between the axial and polar sectors can be dropped for the computation of th QNM frequencies to linear order in spin Pani (2013). After decomposing the perturbation into spherical harmonic basis, the axial and polar perturbations in the Regge-Wheeler gauge to linear order in spin are given by Kojima (1992); Pani (2013)
$\delta g_{\mu\nu}^{\text{axial}}dx^{\mu}dx^{\nu}=\sum_{\ell m}2\Big{\{}h_{0}^{\ell m}(r,t)\Big{[}S_{\theta}^{\ell m}(\theta,\phi)d\theta+S_{\phi}^{\ell m}(\theta,\phi)$
$d\phi\Big{]}dt+h_{1}^{\ell m}(r,t)\Big{[}S_{\theta}^{\ell m}(\theta,\phi)d\theta+S_{\phi}^{\ell m}(\theta,\phi)d\phi\Big{]}dr\Big{\}}\,,$ (10)
$\delta g_{\mu\nu}^{\text{polar}}{dx^{\mu}dx^{\nu}}=\sum_{\ell m}\Big{[}H_{0}^{\ell m}(r,t)dt^{2}+H_{2}^{\ell m}(r,t)dr^{2}$
$+r^{2}K^{\ell m}(r,t)d\Sigma^{2}+2H_{1}^{\ell m}(r,t)dtdr\Big{]}Y^{\ell m}(\theta,\phi)\,,$ (11)
where $S_{\theta}^{\ell m}(\theta,\phi)=-\partial_{\phi}Y_{\ell m}(\theta,\phi)/\sin\theta$, $S_{\phi}^{\ell m}(\theta,\phi)=\sin\theta\partial_{\theta}Y_{\ell m}(\theta,\phi)$, $Y_{\ell m}(\theta,\phi)$ are the tensor spherical harmonics, and $\ell$ and $m$ are the angular and azimuthal numbers of the perturbation. In the following, we will omit the dependence of the radial and angular functions on the angular and azimuthal numbers for brevity of notation. The time dependence can be factorised out as
@@
where the same holds for $h_{1}(r,t)$, $H_{0}(r,t)$, $H_{1}(r,t)$, $H_{2}(r,t)$, $K(r,t)$.
@@ III.1 Reflectivity of compact objects
For generic compact objects, the reflectivity $|\mathcal{R}|$ is related to the membrane parameters (i.e., shear and bulk viscosity of the fluid) through the boundary conditions in Eqs. (II.3) and (II.3) for axial and polar perturbations, respectively. The computation of the reflectivity for arbitrary frequencies is done via numerical integration of the perturbation equations in Eq. (14) and discussed in Sec. III.1.2. Some features of the reflectivity may be extracted analytically and are discussed in the next few paragraphs. We will also discuss the limitations of the linear-in-spin approximation in the form of spurious effects, specifically superradiance, leading to $|\mathcal{R}|>1$ when the spin parameter $\chi$ becomes large enough. This is physically not allowed as superradiance requires the presence of an ergoregion, which is absent at the linear order in spin. Nevertheless, this spuriousness is understood in terms of the limits of validity of the linear-in-spin approximation.
@@ III.1.2 Numerical computation of the reflectivity : Results
Fig. 2 shows the absolute value of the reflectivity of a compact object with $\epsilon\ll 1$ and different values of the shear viscosity, particularly $\eta\in(0,\eta_{\rm{BH}})$ in the top panel and $\eta\in(\eta_{\rm{BH}},+\infty)$ in the bottom panel. We have set $\chi=0.1$ for simplicity. The plots are for axial perturbations, but they look qualitatively similar in the polar case, yielding to the same results. The reflectivity profile coincides with the BH one when $\eta=\eta_{\rm{BH}}$, and describes a perfectly reflecting object in the $\eta\to 0$ and $\eta\to+\infty$ limits. The approximate symmetry in the two panels of Fig. 2 can be understood analytically in the large-frequency limit, where the absolute value of the reflectivity is invariant under the transformation $\eta/\eta_{\rm{BH}}\rightarrow\eta_{\rm BH}/\eta$, as shown in Eq. (16) of Ref. Maggio et al. (2020). An example of this apparent symmetry is provided by the wormhole, which is a purely reflecting object that can be described by either Dirichlet or Neumann boundary conditions, corresponding to $\eta\to 0$ and $\eta\to\infty$ Cardoso et al. (2016a); Maggio et al. (2017, 2019b). While the absolute value of the reflectivity may be similar in the large-frequency limit, the quasinormal mode spectrum is generally different as seen in the differing oscillatory pattern at smaller values of frequency. Note that the oscillatory structure in reflectivity becomes sharply peaked when $\eta\to 0$ and $\eta\to+\infty$. This behaviour is due to the QNMs of the compact object, where the frequencies at which the dips (or resonances) occur coincide with the real part of the QNMs of the object. Using the analytical formulae for the QNMs of ultracompact objects in the small-frequency regime Maggio et al. (2019b, 2021b), we check that the distance between subsequent dips (or resonances) is given by
@@
Let us now analyse the variation of the reflectivity with the object’s compactness for a fixed viscosity. For this purpose, we fix the viscosity in the axial boundary condition in Eq. (II.3) to the BH value. The purely ingoing boundary condition is recovered for $\epsilon\rightarrow 0$, as seen from Eq. (32). However, in the limit $\epsilon\rightarrow 1/2$, corresponding to the membrane radius located at the light ring, $r_{0}\rightarrow 3M$, the reflectivity becomes unity as the axial boundary condition tends to the Dirichlet boundary condition, i.e., $\psi_{\rm{RW}}(r_{0})=0$. This happens regardless of the spin as the second and third terms in the right hand side of Eq. (32) are divergent. The top panel of Fig. 3 shows the variation of the reflectivity with the object’s compactness for axial modes in the slowly spinning case where $\chi=0.1$. The reflectivity coincides with the BH one when $\eta=\eta_{\rm{BH}}$ and $\epsilon$ is small enough, and describes a perfectly reflecting object in the $\epsilon\to 1/2$ limit.
@@
In the polar case, interestingly, the boundary condition in the spinless case ($\chi=0$) does not reach the perfectly reflecting limit as $\epsilon\rightarrow 1/2$. Moreover, the reflectivity depends on the choice of the bulk viscosity of the membrane fluid. However, the linear-in-spin term $H(r_{0},\omega,\eta,\zeta)$ in Eq. (II.3) is divergent in the limit $\epsilon\rightarrow 1/2$, thus changing the behaviour even for a small value of spin. The bottom panel of Fig. 3 shows the variation of the reflectivity with the object’s compactness for polar modes in the slowly spinning case where $\chi=0.1$. Note that the small-spin approximation breaks down for axial and polar modes in the limit $\epsilon\rightarrow 1/2$. This is because the linear-in-spin term of the boundary condition scales as $(\epsilon-1/2)^{-2}$, whereas the spinless term of the boundary condition scales as $(\epsilon-1/2)^{-1}$, as shown in Eq. (II.3). Nevertheless, due to the boundary condition diverging in the limit $\epsilon\rightarrow 1/2$ both in the nonspinning and spinning cases, for $\epsilon=1/2$ the compact object is purely reflecting. It is important to mention that for a given model of compact object, the viscosities ($\eta$, $\zeta$) are uniquely determined and will depend on the compactness parameter $\epsilon$. For simplicity, here we keep the viscosities fixed while varying the compactness (and vice versa). Model-specific studies are left for future work.